### v1

@@ 4 A New Large Neighborhood Search Operator
This paper proposes a new, efficient neighborhood search operator for the Inventory Routing Problem (IRP). This operator is designed to be broadly applicable, regardless of the metaheuristic context. Specifically, given an incumbent solution and a selected retailer $i$, it removes all visits to it over the planning horizon and optimally selects new visit days, delivery quantities, and delivery locations among existing routes or new direct routes with a single visit, while accounting for all relevant costs.
Consequently, this move functions as a simultaneous Reinsertion operator across multiple days, specifically designed for IRP. It will be iteratively applied to all retailers in random order until a local minimum is reached. The closest comparable concept in the literature is the pattern improvement (PI) operator proposed in Vidal et al. (2012), which optimally selects a visit pattern over an extended planning horizon and insertion positions for the multi-period VRP. However, the operator introduced here for the IRP is significantly different, as it handles inventory level constraints, accounts for possible stock-outs, and solves an underlying combinatorial optimization problem to identify the best reinsertion configuration. As a result, it will be referred to as the delivery schedule (DS) operator throughout the remainder of this document. As will be shown, dynamic programming (DP) algorithms can be designed to find optimal reinsertion positions and quantities. The proposed method is versatile and can be adapted to different IRP variants, such as those with penalized stock-outs.
@@
Figure 1 illustrates how one DS move works. One of the retailers (highlighted in blue in the figure) is selected to be removed from its original visit positions, as shown in the first row. After applying the DP algorithm explained below, new insertion positions are determined, as depicted in the second row. Note that the delivery quantities for each day have also been adjusted, leading to completely different inventory levels over the planning horizon. Importantly, while the initial situation is associated with a stock-out quantity equal to $2$ in the third period, the move results in a solution with no stock-outs, as illustrated in the bars close to each plot.
@@ 4.1 Dynamic Programming Algorithm
Functions $C_{t}(I_{i}^{t})$ are piecewise linear, and it is possible to compute them recursively using DP, from $t-1$ to $t$, as illustrated in Figure 2. In practice, these functions will be efficiently represented as linked lists of function pieces, ordered by increasing inventory value, each of which is represented by the endpoints and slope. Moreover, due to Property 1, any segment that does not span an integer inventory value can be discarded throughout the algorithm. Thus, the complexity of the DP algorithm will depend on the number of pieces of this function, as analyzed in Section 4.4.
@@
The cost function $C_{t}(I_{i}^{t})$ is computed in two steps for each $t$. First of all, we compute a delivery cost function $\hat{C}_{t}(\hat{I}^{t}_{i})$ considering inventory levels $\hat{I}^{t}_{i}\in[-d_{i}^{t},U_{i}-d_{i}^{t}]$ that are possibly negative and trigger up to $d_{i}^{t}$ units of stockout at the current period $t$. Then, we include the related stockout penalties to derive $C_{t}(I_{i}^{t})$ for $I_{i}^{t}\in[0,U_{i}-d_{i}^{t}]$. Function $\hat{C}_{t}(\hat{I}^{t}_{i})$, can be computed by taking the minimal cost out of two options:
@@
Finally, we must derive function $C_{t}(I_{i}^{t})$ considering stockout penalties. Note that, thanks to the way stockout costs are defined ($\rho h_{i}$ with $\rho>1$), it is always convenient to avoid any unnecessary stockout leading to $I_{i}^{t}>0$ in any given day. Otherwise, when a stockout occurs (i.e., $\hat{I}_{i}^{t}\leq 0$ in Equation (19)), a penalty of $\rho h_{i}$ multiplied by the number of stockout units ($-\hat{I}_{i}^{t}$) is incurred, and the additional holding cost term $h_{i}\hat{I}_{i}^{t}$ in Equation (19) should be removed accordingly. Therefore, we can calculate the function as:
@@
Observe that, when stockouts are forbidden, we take $C_{t}(I_{i}^{t})=\hat{C}_{t}(\hat{I}_{i}^{t})$ and consider only cases where $\hat{I}_{i}^{t}\geq 0$. Once the DP algorithm has reached day $H$, it determines the inventory level $I_{i}^{H*}$ associated with the minimum cost $C_{H}(I_{i}^{H})$. i.e., $I_{i}^{H*}=\arg\min_{I_{i}^{H}}C_{H}(I_{i}^{H})$. Finally, a backtracking procedure, starting from $I_{i}^{H*}$, determines the quantity $q_{i}^{t}$ to be delivered to the retailer on each day $t$.
@@ 4.2 Preprocessing Functions $F_{t}(q_{i}^{t})$
In alignment with the penalty framework employed to diversify the search in Vidal et al. (2012), Vidal (2022), we allow penalized capacity violations in the routes, weighted by a penalty factor $\omega$. Based on these definitions, we have $F_{t}(q_{i}^{t})=\min_{k\in\mathcal{K}}\left(\gamma_{k}+\omega\max(0,q_{i}^{t}-\kappa_{k})\right)$. This function can be built in $O(K\log(K))$ time by ranking delivery options in increasing order of detour cost, resulting in a list $((\kappa_{1},\gamma_{1}),\dots,(\kappa_{m},\gamma_{m}))$ of $m\leq K$ delivery options, and eliminating dominated options. An illustration of the resulting function is presented in Figure 3.
@@ 4.3 Practical Implementation
  • Infimal convolution: Compute $g(x)=\min_{y}(f_{1}(x-y)+f_{2}(y))$ (see, e.g., Bauschke and Combettes 2017). This operation is more complex and occurs during the calculations of Equation (18). It can be done in $O(\Phi(f_{1})\Phi(f_{2}))$ elementary operations by observing that $g(x)$ is the lower envelope of the infimal convolutions of all pairs of segments in $f_{1}$ and $f_{2}$. Each such convolution can be geometrically calculated as the lower edges of a parallelogram, as illustrated in Figure 4. The resulting function $g$ can have $O(\Phi(f_{1})\times\Phi(f_{2}))$ pieces.
@@ 4.4 Computational Complexity
Preprocessing the functions $F_{t}(q_{i}^{t})$ for all days requires $O(HK\log(K))$ time (Section 4.2). Next, the computational effort to resolve the DP recursion depends on the number of pieces in the $C_{t}(I_{i}^{t})$ functions. Since segments that do not contain integer values in the range $\{0,\dots,U_{i}\}$ can be disregarded, the number of pieces is in $O(U_{i})$. The most computationally intensive operation in each recursion is the calculation of the infimal convolution between the functions $C_{t-1}$ and $F_{t}$, which have $O(U_{i})$ and $O(K)$ pieces, respectively. This convolution requires $O(U_{i}K)$ elementary operations. However, many of these are later discarded—either because they fall outside the valid inventory range or due to domination when retaining the lower-cost segments. As the convolution is performed for each of the $H$ days, exploring the DS neighborhood for a single retailer $i$, including preprocessing, takes $O(U_{i}KH)$ time. This analysis assumes $\log(K)\leq U_{i}$, which typically holds in practice. Moreover, as demonstrated in our computational experiments in Section 6.2, the number of pieces in the function $C_{t}$ remains far below their theoretical maximum in practice. For our problem instances, the complete evaluation of the DS neighborhood for any retailer $i$ is typically completed in a hundred microseconds.
@@ 5 Hybrid Genetic Search
The DS operator described above is integrated into the HGS framework introduced by Vidal et al. (2012), which we adapt to solve the IRP. While the DS operator can be embedded in any metaheuristic, HGS is selected for its versatility and proven effectiveness across numerous VRP variants, including multi-period, multi-depot, and time-window constraints (Vidal et al. 2013, 2014).
@@
After generating an initial population (Section 5.2), the algorithm iteratively applies selection and crossover (Section 5.3), followed by route decoding via the Split algorithm (Prins 2004). A local search phase (Section 5.4) then refines each offspring through (i) route improvement (RI) while keeping delivery quantities fixed, (ii) delivery schedule improvement (DSI) using the DS operator, and (iii) re-optimization of routes (RI). After educating the offspring, the algorithm inserts the resulting solution in the correct subpopulation according to its feasibility and possibly repairs it to additionally include the resulting solution in the feasible subpopulation. If the maximum size of any subpopulation is reached, a survivor selection process is triggered to reduce this subpopulation size (Section 5.5). Subsequently, the penalty parameter associated with capacity violations is adjusted. If the best solution does not improve within a predefined number of iterations, the population is diversified by introducing new solutions generated through the same procedure as in the initialization.
Figure 5: Sketch of the HGS-IRP algorithm.
@@
Each individual in the genetic algorithm represents a solution. It contains, on each day, the sequence of visits to retailers without trip delimiters, i.e., represented as a “giant tour” (Prins 2004). This representation is advantageous since it is possible to optimally split this giant tour to create individual routes using a linear-time split algorithm (Vidal 2016). For the IRP, additional information is needed to completely represent solutions: the quantities delivered to each retailer on each day. Therefore, each solution includes, on each day, (i) a giant tour chromosome encoding the visit sequence and (ii) a quantity chromosome specifying the corresponding delivery quantities.
@@ 5.3 Selection and Crossover
For days with indices strictly below $j_{1}$, a segment of deliveries is inherited from Parent 1. This segment is determined by selecting a random start and end position within the parent’s sequence for that day, as in the PIX crossover of Vidal et al. (2012). The corresponding deliveries, along with their respective delivery quantities, are transferred to the offspring. On days with indices within $\{j_{1},\dots,j_{2}-1\}$, no deliveries are inherited. Finally, for the remaining days, the entire delivery sequence from Parent 1 is copied to the offspring.
@@ 5.4 Education
The local-search education procedure improves solutions by reorganizing the delivery schedule across and routes all days. This operation is applied to complete solutions. The process follows three sequential steps: RI-DSI-RI. The first and last steps, referred to as Route Improvement (RI), involve a standard local search procedure identical to that of Vidal et al. (2012), based on nine classic neighborhood operators, including Relocate and Swap (involving up to two consecutive nodes), as well as 2-Opt and 2-Opt*. These neighborhoods are explored in a randomized order, with any improving move immediately accepted until a local minimum is reached. The intermediate step, Delivery Schedule Improvement (DSI), as described in Section 4, iterates over all retailers to optimize their delivery schedules by reassigning visit days. This phase stops when no further improvements is found in any retailer’s delivery plan.
@@ 5.5 Diversity Management
When a subpopulation reaches its maximum size, the algorithm employs a survivor selection process, similar to the method proposed in Vidal et al. (2012), to determine which individuals are retained for the next generation. This mechanism aims to maintain diversity in visit patterns while preserving elite solutions based on cost. Consequently, discarded individuals are either clones or solutions that perform poorly in terms of both cost and contribution to population diversity.
@@
where $\mathds{1}(\pi_{i}(P_{1})\neq\pi_{i}(P_{2}))$ equals $1$ if the set of delivery days $\pi_{i}(P_{1})$ assigned to retailer $i$ in $P_{1}$ differs from that in $P_{2}$, and 00 otherwise.
As in Vidal et al. (2012), this distance measure is used to define a biased fitness function that influences selection probability, ensuring that the survivor selection process effectively maintains diversity while eliminating redundant or low-quality solutions.
@@ 6 Computational experiments
Benchmark Instances. We test the algorithm on established IRP benchmark instances, categorized into small and large instance datasets. The small dataset consists of well-known instances from (Archetti et al. 2007), while the large instances were introduced in Archetti et al. (2012).
Each instance in the small dataset is defined by three parameters: the planning horizon ($H$), the number of retailers ($n$), and the inventory cost level ($C$). The dataset is divided into four categories, covering two values of $H$ ($H=3$ and $H=6$) and two levels of inventory cost: low inventory cost (LC) and high inventory cost (HC). The number of retailers $n$ in each category ranges from 10 to 50. Each problem size includes five distinct instances. The number of available vehicles $K$ varies between 2 and 5, resulting in a total of 800 instances in the small dataset. The large dataset introduced in Archetti et al. (2012) follows a similar structure, with $H=6$ and $n\in\{50,100,200\}$. The settings for $C$ and $K$ remain consistent with the small instances. For each of the 24 possible parameter combinations ($n$, $H$, $C$, $K$), 10 randomly generated instances were created, leading to a total of 240 large instances.
More recently, Skålnes et al. (2024) introduced a new set of IRP instances designed to be more challenging than the benchmark instances described above. We refer to the small and large datasets as benchmark instances, while we denote the instances from Skålnes et al. (2024) as SAHS instances (named as the initials of the authors). In Section 6.1, we evaluate the HGS algorithm on SAHS instances as well. These instances feature varying numbers of retailers (10 to 200), planning horizons of 6, 9, or 12 periods, and multiple vehicle types with capacities of 8, 18, or 38 EUR-pallets. Unlike the benchmark dataset, which assumes constant retailer demand, SAHS instances incorporate variable demand over time to better simulate real-world conditions.
For the tests in Section 6.3, where stock-out is allowed, we used the small benchmark dataset and set the value of the stock-out penalty $\rho$ as follows: 50, 100, 300, and 1,000,000, the latter being set to a large value to avoid stock-out. Thus, we obtained five instances for each original IRP instance.

### v2

@@ 4 A New Large Neighborhood Search Operator
This paper proposes a new, efficient neighborhood search operator for the Inventory Routing Problem (IRP), designed to be broadly applicable, regardless of the metaheuristic context. Specifically, given an incumbent solution and a selected retailer $i$, it removes all visits to it over the planning horizon and optimally selects new visit days, delivery quantities, and delivery locations among existing routes or new direct routes with a single visit, while accounting for all relevant costs.
Consequently, this move can also be viewed as a simultaneous reinsertion operator across multiple days, specifically designed for IRP. It is iteratively applied to all retailers in random order until a local minimum is reached. The closest comparable concept in the literature is the pattern improvement (PI) operator proposed in vidal2012hybrid, which optimally selects a visit pattern over an extended planning horizon and insertion positions for the multi-period VRP. However, the operator introduced here for the IRP is significantly different, as it handles inventory level constraints, accounts for possible stock-outs, and solves an underlying combinatorial optimization problem to identify the best reinsertion configuration. As a result, it will be referred to as the delivery schedule (DS) operator throughout the remainder of this document. As will be shown, dynamic programming (DP) algorithms can be designed to find optimal reinsertion positions and quantities. The proposed method is versatile and can be adapted to different IRP variants, such as those with penalized stock-outs.
@@
Figure 1 illustrates how the DS operator works. One of the retailers (highlighted in blue in the figure) is selected to be removed from its original visit positions, as shown in the first row. After applying the DP algorithm explained below, new insertion positions are determined, as depicted in the second row. Note that the delivery quantities for each day have also been adjusted, leading to completely different inventory levels over the planning horizon. Importantly, while the initial situation is associated with a stock-out quantity equal to $2$ in the third period, the move results in a solution with no stock-outs, as illustrated in the bars close to each plot.
@@ 4.1 Dynamic Programming Algorithm
Functions $C_{t}(I_{i}^{t})$ are piecewise linear, and it is possible to compute them recursively using DP, from t1t-1 to $t$, as illustrated in Figure 2. In practice, these functions will be efficiently represented as linked lists of function pieces, ordered by increasing inventory value, each of which is represented by the endpoints and slope. Moreover, due to Property 1, any segment that does not span an integer inventory value can be discarded throughout the algorithm. Thus, the complexity of the DP algorithm will depend on the number of pieces of this function, as analyzed in Section 4.4.
@@
The cost function $C_{t}(I_{i}^{t})$ is computed in two steps for each $t$. First of all, we compute a delivery cost function $\hat{C}_{t}(\hat{I}^{t}_{i})$ considering inventory levels $\hat{I}^{t}_{i}\in[-d_{i}^{t},U_{i}-d_{i}^{t}]$ that are possibly negative and trigger up to $d_{i}^{t}$ units of stock-out at the current period $t$. Then, we include the related stock-out penalties to derive $C_{t}(I_{i}^{t})$ for $I_{i}^{t}\in[0,U_{i}-d_{i}^{t}]$. Function $\hat{C}_{t}(\hat{I}^{t}_{i})$, can be computed by taking the minimal cost out of two options:
@@
Finally, we must derive function $C_{t}(I_{i}^{t})$ considering stock-out penalties. Note that, thanks to the way stock-out costs are defined ($\rho h_{i}$ with $\rho>1$), it is always convenient to avoid any unnecessary stock-out leading to $I_{i}^{t}>0$ in any given day. Otherwise, when a stock-out occurs (i.e., $\hat{I}_{i}^{t}\leq 0$ in Equation (19)), a penalty of $\rho h_{i}$ multiplied by the number of stock-out units ($\hat{I}_{i}^{t}$) is incurred, and the additional holding cost term $h_{i}\hat{I}_{i}^{t}$ in Equation (19) should be removed accordingly. Therefore, we can calculate the function as:
@@
Observe that, when stock-outs are forbidden, we can take $C_{t}(I_{i}^{t})=\hat{C}_{t}(\hat{I}_{i}^{t})$ considering only cases where $\hat{I}_{i}^{t}\geq 0$. Moreover, backlogging (allowing negative inventory levels to be carried over and fulfilled in future periods) is not considered in the present paper, but the DS operator could accommodate backorders with minimal structural changes as explained in Appendix 8.
Once the DP algorithm has reached day $H$, it determines the inventory level $I_{i}^{H*}$ associated with the minimum cost $C_{H}(I_{i}^{H})$. i.e., $I_{i}^{H*}=\arg\min_{I_{i}^{H}}C_{H}(I_{i}^{H})$. Finally, a backtracking procedure, starting from $I_{i}^{H*}$, determines the quantity $q_{i}^{t}$ to be delivered to the retailer on each day $t$.
@@ 4.2 Preprocessing Functions $F_{t}(q_{i}^{t})$
In alignment with the penalty framework employed to diversify the search in vidal2012hybrid, vidal2022hybrid, we allow penalized capacity violations in the routes, weighted by a penalty factor $\omega$ which remains fixed during each execution of the DS operator. Based on these definitions, we have $F_{t}(q_{i}^{t})=\min_{k\in\mathcal{K}}\left(\gamma_{k}+\omega\max(0,q_{i}^{t}-\kappa_{k})\right)$. This function can be built in $O(K\log(K))$ time by ranking delivery options in increasing order of detour cost, resulting in a list $((\kappa_{1},\gamma_{1}),\dots,(\kappa_{m},\gamma_{m}))$ of $m\leq K$ delivery options, and eliminating dominated options. An illustration of the resulting function is presented in Figure 3.
@@ 4.3 Practical Implementation
  • Infimal convolution: Compute $g(x)=\min_{y}(f_{1}(x-y)+f_{2}(y))$ (see, e.g., Bauschke2017). This operation is more complex and occurs during the calculations of Equation (18). It can be done in $O(\Phi(f_{1})\Phi(f_{2}))$ elementary operations by observing that $g(x)$ is the lower envelope of the infimal convolutions of all pairs of segments in $f_{1}$ and $f_{2}$. Each such convolution can be geometrically calculated as the lower edges of a parallelogram, as illustrated in Figure 4. The resulting function $g$ can have $O(\Phi(f_{1})\times\Phi(f_{2}))$ pieces.
@@ 4.4 Computational Complexity
Preprocessing the functions $F_{t}(q_{i}^{t})$ for all days requires $O(HK\log(K))$ time (Section 4.2). Next, the computational effort to resolve the DP recursion depends on the number of pieces in the $C_{t}(I_{i}^{t})$ functions. Since segments that do not contain integer values in the range $\{0,\dots,U_{i}\}$ can be disregarded, the number of pieces is in $O(U_{i})$. The most computationally intensive operation in each recursion is the calculation of the infimal convolution between the functions $C_{t-1}$ and $F_{t}$, which have $O(U_{i})$ and $O(K)$ pieces, respectively. This convolution requires $O(U_{i}K)$ elementary operations. However, many of these are later discarded, either because they fall outside the valid inventory range or due to domination when retaining the lower-cost segments. As the convolution is performed for each of the $H$ days, exploring the DS neighborhood for a single retailer $i$, including preprocessing, takes $O(U_{i}KH)$ time. This analysis assumes $\log(K)\leq U_{i}$, which typically holds in practice. Moreover, as demonstrated in our computational experiments in Section 6.2, the number of pieces in the function $C_{t}$ remains far below their theoretical maximum in practice. For our problem instances, the complete evaluation of the DS neighborhood for any retailer $i$ is typically completed in a hundred microseconds.
@@ 5 Hybrid Genetic Search
The DS operator described above is integrated into the HGS framework initially introduced by vidal2012hybrid, which we adapt to solve the IRP. While the DS operator can in principle be embedded within any metaheuristic, we selected HGS due to its flexibility, ability to effectively solve multi-period problems with time constraints, and strong empirical performance across various VRP settings (vidal2013hybrid, vidal2014unified). In addition to its successful application in prior research, HGS has also demonstrated top-tier performance in recent benchmarking competitions, such as the DIMACS VRP Challenge and the EURO/NeurIPS Vehicle Routing Competition (Kool2023), where it formed the core of leading solution methods. These results further support the robustness of HGS as a general-purpose search framework for complex VRPs.
@@
After generating an initial population (Section 5.2), the algorithm iteratively applies selection and crossover (Section 5.3), followed by route decoding via the Split algorithm (prins2004simple). A local search phase (Section 5.4) then refines each offspring through (i) route improvement (RI) while keeping delivery quantities fixed, (ii) delivery schedule improvement (DSI) using the DS operator, and (iii) re-optimization of routes (RI). After educating the offspring, the algorithm inserts the resulting solution in the correct subpopulation according to its feasibility and possibly repairs it to additionally include the resulting solution in the feasible subpopulation. If the maximum size of any subpopulation is reached, a survivor selection process is triggered to reduce this subpopulation size (Section 5.5). Subsequently, the penalty parameter associated with capacity violations is adjusted. If the best solution does not improve within a predefined number of iterations, the population is diversified by introducing new solutions generated through the same procedure as in the initialization.
Figure 5: High-level description of the HGS-IRP algorithm
@@
Each individual in the genetic algorithm represents a solution. It contains, on each day, the sequence of visits to retailers without trip delimiters, i.e., represented as a “giant tour” (prins2004simple). This representation is advantageous since it is possible to optimally split this giant tour to create individual routes using a linear-time split algorithm (Vidal2016). For the IRP, additional information is needed to completely represent solutions: the quantities delivered to each retailer on each day. Therefore, each solution includes, on each day, (i) a giant tour chromosome encoding the visit sequence and (ii) a quantity chromosome specifying the corresponding delivery quantities.
@@ 5.3 Selection and Crossover
For days with indices strictly below $j_{1}$, a segment of deliveries is inherited from Parent 1. This segment is determined by selecting a random start and end position within the parent’s sequence for that day, as in the PIX crossover of vidal2012hybrid. The corresponding deliveries, along with their respective delivery quantities, are transferred to the offspring. On days with indices within $\{j_{1},\dots,j_{2}-1\}$, no deliveries are inherited. Finally, for the remaining days, the entire delivery sequence from Parent 1 is copied to the offspring.
@@ 5.4 Education
The local-search education procedure improves solutions by reorganizing the delivery schedule across all routes and days. This operation is applied to complete solutions. The process follows three sequential steps: RI-DSI-RI. The first and last steps, referred to as Route Improvement (RI), involve a standard local search procedure identical to that of vidal2012hybrid, based on nine classic neighborhood operators, including Relocate and Swap (involving up to two consecutive nodes), as well as 2-Opt and 2-Opt*. These neighborhoods are explored in a randomized order, with any improving move immediately accepted until a local minimum is reached. For the Delivery Schedule Improvement (DSI) phase, as described in Section 4, we iterate over all retailers in random order to optimize their delivery schedules by reassigning visit days. This phase stops when a complete pass over all retailers results in no improvement in any retailer’s delivery plan.
@@ 5.5 Diversity Management
When a subpopulation reaches its maximum size, the algorithm employs a survivor selection process, similar to the method proposed in vidal2012hybrid, to determine which individuals are retained for the next generation. This mechanism aims to maintain diversity in visit patterns while preserving elite solutions based on cost. Consequently, discarded individuals are either clones or solutions that perform poorly in terms of both cost and contribution to population diversity.
@@
where $\mathds{1}(\pi_{i}(P_{1})\neq\pi_{i}(P_{2}))$ equals $1$ if the set of delivery days $\pi_{i}(P_{1})$ assigned to retailer $i$ in $P_{1}$ differs from that in $P_{2}$, and 0 otherwise.
As in vidal2012hybrid, this distance measure is used to define a biased fitness function that influences selection probability, ensuring that the survivor selection process effectively maintains diversity while eliminating redundant or low-quality solutions.
@@ 6 Computational experiments
Benchmark Instances. We test the algorithm on established IRP benchmark instances, categorized into small and large instance datasets. The small dataset consists of well-known instances from (archetti2007branch), while the large instances were introduced in archetti2012hybrid.
Each instance in the small dataset is defined by three parameters: the planning horizon ($H$), the number of retailers ($n$), and the inventory cost level ($C$). The dataset is divided into four categories, covering two values of $H$ ($H=3$ and $H=6$) and two levels of inventory cost: low inventory cost (LC) and high inventory cost (HC). The number of retailers $n$ in each category ranges from 10 to 50. Each problem size includes five distinct instances. The number of available vehicles $K$ varies between 2 and 5, resulting in a total of 800 instances in the small dataset. The large dataset introduced in archetti2012hybrid follows a similar structure, with $H=6$ and $n\in\{50,100,200\}$. The settings for $C$ and $K$ remain consistent with the small instances. For each of the 24 possible parameter combinations ($n$, $H$, $C$, $K$), 10 randomly generated instances were created, leading to a total of 240 large instances. More recently, skaalnes2024new introduced a new set of IRP instances designed to be more challenging than the benchmark instances described above. We refer to the small and large datasets as benchmark instances, while we denote the instances from skaalnes2024new as SAHS instances (named as the initials of the authors). In Section 6.1, we evaluate the HGS algorithm on SAHS instances as well. These instances feature varying numbers of retailers (10 to 200), planning horizons of 6, 9, or 12 periods, and multiple vehicle types with capacities of 8, 18, or 38 EUR-pallets. Unlike the benchmark dataset, which assumes constant retailer demand, SAHS instances incorporate variable demand over time to better simulate real-world conditions.
For the tests in Section 6.3, where stock-out is allowed, we used the small benchmark dataset and set the value of the stock-out penalty $\rho$ as follows: 50, 100, 300, and 1,000,000, the latter being set to a large value to avoid stock-out. Thus, we obtained four instances for each original IRP instance.